OpenStudy (anonymous):

Find an equation of the parabola having vertex (-1, 4) and containing the point (1, 2).

OpenStudy (anonymous):

a. y+4=-1/2(x-1)^2 b. y+4=-1/2(x-1)^2 c. y-4=-1/2(x+1)^2